OpenStudy (anonymous):

Jeff stores coffee powder in a steel container with a capacity of 252.3 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 6.2 inches, what is its radius? Answer 1.8 inches 2.1 inches 3.1 inches 3.6 inches

OpenStudy (dumbcow):

V = pi*h*r^2 252.3 = 6.2pi *r^2 r = sqrt(252.3/6.2pi)

OpenStudy (anonymous):

so would the answer be 3.1 ? im so confused

OpenStudy (dumbcow):

that expression is the radius, put it in a calculator to find out the decimal answer

OpenStudy (anonymous):

so i divide 252.3 by 6.2?

OpenStudy (dumbcow):

252.3/(6.2*pi)

OpenStudy (anonymous):

and i got 12.9

OpenStudy (dumbcow):

what should you do next?

OpenStudy (anonymous):

divide it by 2?

OpenStudy (anonymous):

wait no thats wrong.. i have no idea

OpenStudy (anonymous):

?